OpenStudy (babyslapmafro):

Please help me we the following problem dx/dt=d/dt [t^2+4/7t]

OpenStudy (babyslapmafro):

=d/dt [(t^2/7t)+(4/7t)]

OpenStudy (babyslapmafro):

=d/dt [(t/7)+(4/7t)]

OpenStudy (babyslapmafro):

=d/dt [t/7] + d/dt [4/7t]

OpenStudy (babyslapmafro):

now i'm stuck

OpenStudy (anonymous):

=d/dt [t/7] + d/dt [4t^(-1)/7] =1/7 - 4/7 *(-1) t^(-1-1)

OpenStudy (anonymous):

So, dx/dt= 1/7 +4/7t^2

OpenStudy (anonymous):

Does it help?

OpenStudy (babyslapmafro):

d/dt [4t^(-1)/7] is there a rule that explains how you got that?

OpenStudy (babyslapmafro):

what would happen if it were d/dt [4t/7] ?

OpenStudy (anonymous):

Sorry I was out

OpenStudy (anonymous):

Yes there is a rule.... d/dt [4t/7] =4/7

OpenStudy (anonymous):

Now, This may help: d(ax^n)/dx = a*n x^(n-1)

OpenStudy (amistre64):

\[x'=(t^2+4/7t)'\] \[x'=(t^2)'+(4/7t)'\] \[x'=(2t)+(4/7t)'\] where is the "t" in that last term, top or bottom?

OpenStudy (amistre64):

for that matter, i may have interpreted the setup incorrectly. \[\frac{t^2+4}{7t}~??\] \[t^2+\frac{4}{7t}~??\] \[t^2+\frac{4}{7}t~??\]